# STIQ: Safeguarding Training and Inferencing of Quantum Neural Networks from Untrusted Cloud

### v1

@@ Abstract.
The high expenses imposed by current quantum cloud providers, coupled with the escalating need for quantum resources, may incentivize the emergence of cheaper cloud-based quantum services from potentially untrusted providers. Deploying or hosting quantum models, such as Quantum Neural Networks (QNNs), on these untrusted platforms introduces a myriad of security concerns, with the most critical one being model theft. This vulnerability stems from the cloud provider’s full access to these circuits during training and/or inference. In this work, we introduce STIQ, a novel ensemble-based strategy designed to safeguard QNNs against such cloud-based adversaries. Our method innovatively trains two distinct QNNs concurrently, hosting them on same or different platforms, in a manner that each network yields obfuscated outputs rendering the individual QNNs ineffective for adversaries operating within cloud environments. However, when these outputs are combined locally (using an aggregate function), they reveal the correct result. Through extensive experiments across various QNNs and datasets, our technique has proven to effectively masks the accuracy and losses of the individually hosted models by upto 76%, albeit at the expense of 2$\leq$ 2$\times$ 2 × increase in the total computational overhead. This trade-off, however, is a small price to pay for the enhanced security and integrity of QNNs in a cloud-based environment prone to untrusted adversaries. We also demonstrated STIQ’s practical application by evaluating it on real 127-qubit IBM_Sherbrooke hardware, showing that STIQ achieves up to 60% obfuscation, with combined performance comparable to an unobfuscated model.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Quantum computing is rapidly progressing, with companies like Atom Computing and IBM recently unveiling the largest quantum processors ever developed, boasting 1,225 and 1,121 qubits, respectively (computing2023quantum, ; gambetta2023hardware, ). The significant interest in quantum computing among academic and research communities stems from its potential to offer substantial computational speedups over classical computers for certain problems. Researchers have already begun leveraging these noisy intermediate-scale quantum (NISQ) machines to demonstrate practical utility in this pre-fault-tolerant era (kim2023evidence, ). Within this emergent field, quantum machine learning (QML) has also gained considerable attention, merging the power of quantum computing with classical machine learning algorithms. QML heuristically explores the potential of improving learning algorithms by leveraging the unique capabilities of quantum computers, opening new horizons in computational speed and capability. Several QML models have been explored, including quantum support vector machines (QSVMs) (rebentrost2014quantum, ), quantum generative adversarial networks (QGANs) (dallaire2018quantum, ), and quantum convolutional neural networks (QCNNs) (cong2019quantum, ). However, quantum neural networks (QNNs) (schuld2014quest, ; abbas2021power, ) stand out as the most notable development, mirroring the structure and function of classical neural networks within a quantum framework.
@@
Figure 1. Figure depicting the workflow of QMLaaS, showcasing how users can access QNNs through API queries. Deployment of such QNNs on untrusted clouds exposes them to various potential adversarial attacks.
@@
At the core of QNNs lies the parameterized quantum circuit (PQC), which consists of several trainable single-qubit and two-qubit gates. The parameters of these gates are updated using classical optimizers to solve the required task. PQCs have been successfully employed in various algorithms to tackle a wide range of problems, including combinatorial optimization (farhi2014quantum, ), molecular energy calculation (kandala2017hardware, ), classification (farhi2018classification, ), finance (nakaji2022approximate, ) and security (beaudoin2022quantum, ; kundu2022security, ). However, optimizing PQCs, especially in the context of QNNs, presents significant challenges in terms of time, complexity, and quantum resource requirements. This is because, no-cloning theorem, a fundamental principle of quantum mechanics, prevents the duplication of quantum states, making traditional backpropagation techniques used in classical neural networks impractical. To circumvent this limitation, alternative methods like the parameter shift rule are employed which requires executing the quantum circuit multiple times to estimate the gradients needed for optimization. Specifically, each parameter in the circuit needs to be run twice for gradient computation, significantly increasing the quantum resource demand as the number of parameters grows. Thus, as QML expands, we will soon see a rising trend towards hosting uniquely architected QML models or those trained on novel quantum datasets on the cloud (Fig. 1), with paid access leading to the rise of Quantum Machine-Learning-as-a-Service (QMLaaS).
@@
However, as access to current quantum hardware remains limited and extremely expensive, it paves the way for the emergence of third-party quantum cloud providers, which could improve access to quantum computers, but potentially at the cost of compromised security. As a result, even though QMLaaS could democratize access to cutting-edge PQC based models, it also exposes QML models to several adversarial attacks A particularly pressing concern is the risk of ”model theft” attacks. In these scenarios, attackers can steal a model’s architecture and behavior because hosting models on untrusted cloud platforms can provide adversaries with white-box access. This means they have complete insight into the specifics of the QNN circuit like the trained parameters, architecture, etc. Consequently, when a QNN is sent to the cloud for training or hosting, the cloud provider might potentially access and replicate the model. This not only jeopardizes the uniqueness of the QNN but also poses a significant risk of intellectual property theft. The provider could potentially exploit this access to offer a similar service, leveraging the stolen model. Given these risks, there is a need for robust security measures when employing untrusted cloud services for QNNs to prevent against exploitation and safeguard intellectual property.
@@
Several studies have addressed protecting quantum circuits from untrusted compilers (saki2021split, ; suresh2021short, ) or clouds (upadhyay2022robust, ). However, there’s a notable gap in research regarding the protection of QNNs from cloud-based adversaries during training and inference stages. Techniques designed for untrusted compilers are inadequate for cloud scenarios, as they only secure the circuit pre-compilation, allowing cloud-based adversaries full access thereafter. Moreover, while quantum homomorphic encryption (QHE) (dulek2016quantum, ; fisher2014quantum, ) offers theoretical security, it remains impractical due to its significant overhead. Blind quantum computing (BQC) (broadbent2009universal, ; barz2012demonstration, ) requires clients to perform specific operations on a quantum hardware locally, which is often unfeasible. Authors in (patel2023toward, ) suggested adding random X gates to quantum circuits to obfuscate outputs, but this method has vulnerabilities. Determined adversaries could reverse-engineer the circuit by testing various combinations, and patterns in output distribution might reveal the placement of X gates. This situation underscores the urgent need to develop more effective and practical protocols to safeguard QNNs from untrusted cloud providers.
@@
Here, we introduce STIQ (Safeguarding Training and Inferencing of QNNs), a novel approach designed to safeguard QML models against model theft by untrustworthy cloud providers. Our technique draws inspiration from the principles of ensemble learning but differs fundamentally in the training methodology. Unlike traditional ensemble learning, where each model is independently trained to achieve accurate predictions before their outputs are merged to produce the final result, STIQ trains models in parallel, ensuring that each generates obfuscated results with deliberately reduced individual performance. However, these outputs, when aggregated, reconstruct the accurate final prediction. This method ensures that even though untrusted cloud providers may access the parameters and architecture of the QNNs, the data remains protected, as the outputs of individual QNNs are erroneous and misleading to potential adversaries. Our evaluation of STIQ spans various datasets and hardware configurations, showcasing its simplicity and effectiveness in defending QNNs against potential threats posed by cloud-based adversaries. We also tested STIQ on real quantum hardware (IBM_Sherbrooke) and demonstrated the effectiveness of our approach in a more realistic scenario.
To the best of our knowledge, this is the first work that presents an effective solution in safeguarding QNNs against model theft attacks in a cloud-based environment.
2. Background
2.1. Basics of Quantum Computing
@@
2.2. Quantum Neural Network (QNN)
@@
Figure 2. Architecture of a 4-qubit hybrid QNN. Classical features are encoded as angles of quantum rotation gates ($R_{Z}$). PQC transforms encoded states to explore the search space and entangle features. Measured expectation values are then fed into a classical linear layer for final prediction.
@@
3. Related Works and Limitations
@@
Model theft poses a substantial security risk to emerging QML platforms particularly with the advent of third-party quantum cloud providers, where even QNN circuits submitted for training and/or inference are vulnerable. To address the generic quantum circuit security risks, strategies have been explored, such as by authors in (upadhyay2022robust, ), which suggests dividing circuit execution between trusted and untrusted providers to mitigate adversarial interference risks. However, this approach proves ineffective for QNNs, as access by any single untrusted provider could lead to model theft. Cryptography-based techniques like quantum homomorphic encryption (QHE) offers a theoretical solution by enabling computation on encrypted quantum data, thereby safeguarding the privacy of the computation outcomes (yu2014limitations, ; dulek2016quantum, ; fisher2014quantum, ). However, the exponential increase in computational overhead, render it impractical in the near term. Likewise, blind quantum computing (BQC) (broadbent2009universal, ; barz2012demonstration, ; mantri2013optimal, ) allows a user to perform quantum computations on a remote quantum computer without revealing the nature of the computation to the server. However, it still requires the client to have the capability to prepare specific quantum bits and perform certain quantum operations which poses a barrier for several users.
@@
There have also been obfuscation based strategies to protect quantum circuits from less trustworthy cloud providers (patel2023toward, ; upadhyay2022robust, ; kundu2024evaluating, ) and compilers (suresh2021short, ). Despite these efforts, to the best of our knowledge, there have not yet been any methods developed that effectively protect QNNs against model theft. This is a significant concern because circuit obfuscation techniques, while potentially useful in some cases, cannot be straightforwardly applied to QNNs due to several inherent drawbacks. Specifically, in (patel2023toward, ) authors proposed randomly adding X gates at the end of the quantum circuit to obfuscating outputs from adversaries. However, an adversary with sufficient time and resources can systematically brute-force the outputs. Furthermore, patterns in the output distribution or correlations between different runs of the circuit could provide hints about the underlying placement of X gates. Other challenges lies in the training phase of an obfuscated circuit. Specifically, the process of gradient calculation—which is crucial for optimizing the QNN—becomes problematic.
Techniques such as the parameter-shift rule and simultaneous perturbation stochastic approximation (SPSA), commonly employed for gradient calculation in practical quantum computing scenarios, may yield highly erroneous or even irrelevant results when applied to circuits that have been altered through obfuscation. This is because the effectiveness of these gradient calculation techniques is closely tied to the circuit’s specific gate configurations, which are deliberately altered in the process of obfuscation to protect against theft. Furthermore, the performance of obfuscated circuits, when executed on real quantum hardware, tends to suffer, even if by a small amount, due to the inclusion of additional gates. This is because each gate added to a circuit increases the overall gate count and circuit depth, leading to a greater accumulation of gate errors and decoherence errors over the training process. These errors pose major challenges in quantum computing, as they directly contribute to the degradation of model accuracy and efficiency. These dilemmas underscores the challenge in safeguarding QNN circuits against model theft without compromising their operational integrity.
4. Proposed Technique and Implementation
@@
4.1. QNN Assets and Adversary Motivation
@@ 4.2. Threat Model
Figure 3. Figure illustrating the training workflow of STIQ. Both QNN-1 and QNN-2 are trained using a unified loss function that optimally balances classification loss $C()$ and divergence loss $D()$.
4.3. Methodology Overview
@@ 4.4. Training Methodology
Figure 4. Figure depicting the STIQ inference methodology: (a) A user query $X$ is processed by two cloud-hosted QNNs, resulting in obfuscated outputs $y_{1}$ and $y_{2}$. Locally, STIQ aggregates these outputs to produce the correct result $y$ for the user. (b) Illustration of an example explaining how individual models may yield erroneous predictions, yet their combined output through STIQ accurately produces the correct output vector.
@@ 4.5. Inference Methodology
Figure 5. Plot demonstrating the test accuracies of different 4-qubit QNNs trained on the Fashion-4 dataset with varying penalties ($\lambda$). It is clear that while increasing the penalty enhances obfuscation, it also adversely affects the combined performance.
5. Evaluation
5.1. Setup
@@
Device: In our study, we conducted all noiseless experiments on Pennylane’s (bergholm2018pennylane, ) “lightning.qubit” device using the ”Adjoint” differentiation method to calculate the gradients. This method is known for its efficiency in simulating quantum circuits without noise. For our noisy inferencing, we used “qiskit.aer” device. Given the long queue times and limited availability of real quantum devices and our goal to closely replicate the conditions of these noisy quantum devices, we opted to directly integrate noise models from actual quantum hardware. Specifically, we employed noise models from IBM_Sherbrooke, IBM_Cairo and IBM_Hanoi and incorporated them to the ”qiskit.aer” device to simulate the noise characteristics of these machines accurately. Considering that traditional methods for calculating gradients, such as backpropagation, are not viable on actual quantum devices, we utilized the Simultaneous Perturbation Stochastic Approximation (SPSA) method for gradient calculation in our QNN training. This approach produces noisy gradients, which are more indicative of the performance one would expect from real quantum hardware, thereby ensuring our experimental results are more aligned with practical quantum computing environments. We also conducted experiments on real quantum hardware, specifically IBM_Sherbrooke.
Training: For evaluation, we used PQC-4 and 19 from (sim2019expressibility, ) to build our QNNs (PQC-$x$ represents circuit-$x$ in (sim2019expressibility, )), initialized with random weights. In most of the experiments we used a 4-qubit QNN as our cloud-based QNN (if not mentioned otherwise). For embedding classical features to their corresponding quantum state we use the angle encoding technique with $RZ$ gate and for measurement, we calculate the Pauli-Z basis expectation value over all the qubits. For training the QNNs we used CrossEntropyLoss() as our classification loss function and CosineSimilarity() as the divergence loss function. The hyperparameters used for training are; Epochs: 30, learning_rate ($\eta$): 0.01, batch_size = 32 and optimizer: Adam. All training are done on an Intel Core-i7-12700H CPU with 40GB of RAM.
Dataset: Since NISQ devices struggle with large images due to limited qubits, high error rates and complex data encoding, similar to recent works (wang2022qoc, ; fu2024quantumleak, ), we conduct all experiments using a reduced feature set of MNIST, Fashion, Kuzushiji and Letters datasets with latent dimension $d=8$ (from original 28$\times$28 image) generated using a convolutional autoencoder (alam2021quantum, ). Thus, for each dataset, we create a smaller 4-class dataset from these reduced feature sets i.e., MNIST-4 (class 0, 1, 2, 3), Fashion-4 (class 6, 7, 8, 9), Kuzushiji-4 (class 3, 5, 6, 9) and Letters-4 (class 1, 2, 3, 4) with each having 1000 samples (700 for training and 300 for testing). Since we use 4-qubit QNN models for training and each of these datasets is of dimension $d=8$, we encode 2 features per qubit. Number of shots/trials is set to 1000 for all experiments.
5.2. Simulation Results
Penalty ($\lambda$) Analysis: To achieve the goal of balancing obfuscation and performance in our QNN setup, we first need to determine the optimal value for the penalty parameter, $\lambda$, as introduced in Eq. 1. This parameter is critical for adjusting the impact of the divergence loss $L_{D}$, in this case, CosineSimilarity, on the overall training process. Our objective is to find a $\lambda$ value that allows individual models to maintain a high degree of obfuscation without significantly compromising the combined performance of the models when compared to an unobfuscated baseline model. An analysis of this is illustrated in Fig. 5 and Table 1, which shows the test accuracy of a 4-qubit QNN model trained on the Fashion-4 dataset. The figure reveals that as $\lambda$ increases, enhancing the level of obfuscation by giving more weight to the divergence loss, the performance of the individual models, QNN-1 and QNN-2, begins to decline. This decrease in performance suggests that a high penalty on divergence loss can negatively affect the model’s ability to optimize effectively, as the divergence loss begins to dominate over the classification loss. Based on the trends observed in the figure, it becomes clear that an optimal range for $\lambda$ exists. Specifically, for the CosineSimilarity divergence function, a $\lambda$ value within the range of $[0.1,0.3]$ seems to be the optimal range. Within this range, the models achieve a desirable level of obfuscation (45-75%) while still performing closely to the unobfuscated baseline model in terms of accuracy ($\leq 1\%$).
@@ 5.2. Simulation Results
Table 1. Comparison of final test accuracy after training a 4-qubit QNN on Fashion-4 with varying penalties.
@@
Divergence Loss ($L_{D}$): Next, we evaluate the effects of employing various divergence loss functions on the training and performance of 4-qubit QNNs on the MNIST-4 dataset. The divergence loss functions under consideration are L1 loss, L2 loss, and CosineSimilarity. The goal is to understand how each loss function influences the model’s test accuracy, as depicted in Fig. 6 and Table 2. To ensure effective training of the models with each type of loss function, it was necessary to empirically determine suitable penalty values for each case. This step was crucial because the choice of divergence loss function and its corresponding penalty value directly affects the model’s ability to learn. The analysis revealed a notable pattern: while initially, the models trained with L1 and L2 loss functions showed promising performance, their combined performance began to decline after reaching a certain number of epochs. This decline is is because of the divergence loss ($L_{D}$), which, beyond a certain point, starts to divert more from the overall learning process than the classification loss ($L_{C}$) contributes. Essentially, the models become too focused on minimizing $L_{D}$ at the expense of $L_{C}$, leading to poorer overall performance. In contrast, models utilizing CosineSimilarity as the divergence loss function exhibited a more stable training progression. This stability suggests that CosineSimilarity, unlike L1 or L2 loss, provides a balance that prevents the divergence loss from overwhelming the classification loss. Consequently, our findings indicate that CosineSimilarity is a preferable choice for the divergence loss function in the context of training QNNs, as it supports more consistent and effective learning.
Table 2. Final test accuracies after training a QNN on MNIST-4 with different divergence loss functions.
@@
Figure 6. Test accuracies of 4-qubit QNNs trained on the MNIST-4 dataset with various divergence loss functions ($L_{D}$).
Table 3. Comparison of final test accuracy and loss after training a 8-qubit 2-layer QNN on different datasets for 30 epochs with learning rate $0.01$ and penalty ($\lambda$): 0.3. We can see that eventhough performance on individual QNNs are greatly obfuscated, performance of their combined outputs is similar to that of the baseline model.

### v2

@@ Abstract
The high expenses imposed by current quantum cloud providers, coupled with the escalating need for quantum resources, may incentivize the emergence of cheaper cloud-based quantum services from potentially untrusted providers. Deploying or hosting quantum models, such as Quantum Neural Networks (QNNs), on these untrusted platforms introduces a myriad of security concerns, with the most critical one being model theft. This vulnerability stems from the cloud provider’s full access to these circuits during training and/or inference. In this work, we introduce STIQ, a novel ensemble-based strategy designed to safeguard QNNs against such cloud-based adversaries. Our method innovatively trains two distinct QNNs concurrently, hosting them on same or different platforms, in a manner that each network yields obfuscated outputs rendering the individual QNNs ineffective for adversaries operating within cloud environments. However, when these outputs are combined locally (using an aggregate function), they reveal the correct result. Through extensive experiments across various QNNs and datasets, our technique has proven to effectively masks the accuracy and losses of the individually hosted models by upto 76%, albeit at the expense of 2$\leq$ 2$\times$ 2 × increase in the total computational overhead. This trade-off, however, is a small price to pay for the enhanced security and integrity of QNNs in a cloud-based environment prone to untrusted adversaries. We also demonstrated STIQ’s practical application by evaluating it on multiple real quantum hardwares, showing that STIQ achieves up to $\approx$ 70% obfuscation, with combined performance similar to an unobfuscated model.
Index Terms:
obfuscate, quantum neural network, adversary
I Introduction
@@
Quantum computing is rapidly progressing, with companies like Atom Computing and IBM recently unveiling the largest quantum processors ever developed, boasting 1,225 and 1,121 qubits, respectively [1, 2]. The significant interest in quantum computing among academic and research communities stems from its potential to offer substantial computational speedups over classical computers for certain problems. Researchers have already begun leveraging these noisy intermediate-scale quantum (NISQ) machines to demonstrate practical utility in this pre-fault-tolerant era [3]. Within this emergent field, quantum machine learning (QML) has also gained considerable attention, merging the power of quantum computing with classical machine learning algorithms. QML heuristically explores the potential of improving learning algorithms by leveraging the unique capabilities of quantum computers, opening new horizons in computational speed and capability. Several QML models have been explored, including quantum support vector machines (QSVMs) [4], quantum generative adversarial networks (QGANs) [5], and quantum convolutional neural networks (QCNNs) [6]. However, quantum neural networks (QNNs) [7, 8] stand out as the most notable development, mirroring the structure and function of classical neural networks within a quantum framework.
@@
Figure 1: Figure depicting the workflow of QMLaaS, showcasing how users can access QNNs through API queries. Deployment of such QNNs on untrusted clouds exposes them to various potential adversarial attacks.
@@
At the core of QNNs lies the parameterized quantum circuit (PQC), which consists of several trainable single-qubit and two-qubit gates. The parameters of these gates are updated using classical optimizers to solve the required task. PQCs have been successfully employed in various algorithms to tackle a wide range of problems, including combinatorial optimization [9], molecular energy calculation [10], classification [11], finance [12] and security [13, 14]. However, optimizing PQCs, especially in the context of QNNs, presents significant challenges in terms of time, complexity, and quantum resource requirements. This is because, no-cloning theorem, a fundamental principle of quantum mechanics, prevents the duplication of quantum states, making traditional backpropagation techniques used in classical neural networks impractical. To circumvent this limitation, alternative methods like the parameter shift rule are employed which requires executing the quantum circuit multiple times to estimate the gradients needed for optimization. Specifically, each parameter in the circuit needs to be run twice for gradient computation, significantly increasing the quantum resource demand as the number of parameters grows. Thus, as QML expands, we will soon see a rising trend towards hosting uniquely architected QML models or those trained on novel quantum datasets on the cloud (Fig. 1), with paid access leading to the rise of Quantum Machine-Learning-as-a-Service (QMLaaS).
I-A Motivation
@@
As access to current quantum hardware remains limited and extremely expensive, it paves the way for the emergence of third-party quantum cloud providers, which could improve access to quantum computers, but potentially at the cost of compromised security. As a result, even though QMLaaS could democratize access to cutting-edge PQC based models, it also exposes QML models to several adversarial attacks A particularly pressing concern is the risk of ”model theft” attacks. In these scenarios, attackers can steal a model’s architecture and behavior because hosting models on untrusted cloud platforms can provide adversaries with white-box access. This means they have complete insight into the specifics of the QNN circuit like the trained parameters, architecture, etc. Consequently, when a QNN is sent to the cloud for training or hosting, the cloud provider might potentially access and replicate the model. This not only jeopardizes the uniqueness of the QNN but also poses a significant risk of intellectual property theft. The provider could potentially exploit this access to offer a similar service, leveraging the stolen model. Given these risks, there is a need for robust security measures when employing untrusted cloud services for QNNs to prevent against exploitation and safeguard intellectual property.
@@
Several studies have addressed protecting quantum circuits from untrusted compilers [15, 16] or clouds [17]. However, there’s a notable gap in research regarding the protection of QNNs from cloud-based adversaries during training and inference stages. Techniques designed for untrusted compilers are inadequate for cloud scenarios, as they only secure the circuit pre-compilation, allowing cloud-based adversaries full access thereafter. Moreover, while quantum homomorphic encryption (QHE) [18, 19] offers theoretical security, it remains impractical due to its significant overhead. Blind quantum computing (BQC) [20, 21] requires clients to perform specific operations on a quantum hardware locally, which is often unfeasible. Authors in [22] suggested adding random X gates to quantum circuits to obfuscate outputs, but this method has vulnerabilities. Determined adversaries could reverse-engineer the circuit by testing various combinations, and patterns in output distribution might reveal the placement of X gates. Classical techniques for protecting Deep Neural Networks (DNNs) from untrusted cloud environments [23, 24, 25] typically employ matrix transformation or encryption-based strategies. However, these methods are not directly transferrable to QNNs since, the measurement process in QNNs is irreversible, which complicates the application of classical defense strategies that rely on reversible transformations. Furthermore, the quantum counterparts of these encryption techniques tend to introduce significant overhead, making them impractical. This underscores the urgent need to develop more effective and practical protocols to safeguard QNNs from untrusted cloud providers.
I-B Contribution
@@
Here, we introduce STIQ (Safeguarding Training and Inferencing of QNNs), a novel approach designed to safeguard QML models against model theft by untrustworthy cloud providers. Our technique draws inspiration from the principles of ensemble learning but differs fundamentally in the training methodology. Unlike traditional ensemble learning, where each model is independently trained to achieve accurate predictions before their outputs are merged to produce the final result, STIQ trains models in parallel, ensuring that each generates obfuscated results with deliberately reduced individual performance. However, these outputs, when aggregated, reconstruct the accurate final prediction. This method ensures that even though untrusted cloud providers may access the parameters and architecture of the QNNs, the data remains protected, as the outputs of individual QNNs are erroneous and misleading to potential adversaries. Our evaluation of STIQ spans various datasets and hardware configurations, showcasing its simplicity and effectiveness in defending QNNs against potential threats posed by cloud-based adversaries. We also tested STIQ on real quantum hardwares (IBM_Sherbrooke, IBM_Brisbane and IBM_Kyoto) and demonstrated the effectiveness of our approach in a more realistic scenario.
II Background
II-A Basics of Quantum Computing
@@
II-B Quantum Neural Network (QNN)
@@
Figure 2: Architecture of a 4-qubit hybrid QNN. Classical features are encoded as angles of quantum rotation gates ($R_{Z}$). PQC transforms encoded states to explore the search space and entangle features. Measured expectation values are then fed into a classical linear layer for final prediction.
@@
III Related Works and Limitations
@@
Model theft poses a substantial security risk to emerging QML platforms particularly with the advent of third-party quantum cloud providers, where even QNN circuits submitted for training and/or inference are vulnerable. To address the generic quantum circuit security risks, strategies have been explored, such as by authors in [17], which suggests dividing circuit execution between trusted and untrusted providers to mitigate adversarial interference risks. However, this approach proves ineffective for QNNs, as access by any single untrusted provider could lead to model theft. Cryptography-based techniques like quantum homomorphic encryption (QHE) offers a theoretical solution by enabling computation on encrypted quantum data, thereby safeguarding the privacy of the computation outcomes [26, 18, 19]. However, the exponential increase in computational overhead, render it impractical in the near term. Likewise, blind quantum computing (BQC) [20, 21, 27] allows a user to perform quantum computations on a remote quantum computer without revealing the nature of the computation to the server. However, it still requires the client to have the capability to prepare specific quantum bits and perform certain quantum operations which poses a barrier for several users.
@@
There have also been obfuscation based strategies to protect quantum circuits from less trustworthy cloud providers [22, 17, 28] and compilers [16]. Despite these efforts, to the best of our knowledge, there have not yet been any methods developed that effectively protect QNNs against model theft. This is a significant concern because circuit obfuscation techniques, while potentially useful in some cases, cannot be straightforwardly applied to QNNs due to several inherent drawbacks. Specifically, in [22] authors proposed randomly adding X gates at the end of the quantum circuit to obfuscating outputs from adversaries. However, an adversary with sufficient time and resources can systematically brute-force the outputs. Furthermore, patterns in the output distribution or correlations between different runs of the circuit could provide hints about the underlying placement of X gates.
Recent works in the classical domain have addressed secure inference of neural networks on untrusted clouds. For instance, the authors in [23] discuss a method involving the multiplication of neural network data and weight matrices by pseudorandom sparse matrices. This manipulation allows secure execution on the cloud, with subsequent local computations retrieving the correct results. However, this technique cannot be directly applied to QNNs due to the irreversible nature of the measurement process in quantum computing, which complicates the retrieval of correct outputs using classical computations. Additionally, while there have been efforts to implement encryption-based methods [25, 24, 29], these approaches are challenging to integrate into quantum circuits. Those that can be adapted incur significant overheads. These dilemmas highlights the challenge in safeguarding QNN circuits against model theft without compromising their operational integrity.
IV Proposed Technique and Implementation
@@
IV-A QNN Assets and Adversary Motivation
@@ IV-B Threat Model
Figure 3: Figure illustrating the training workflow of STIQ. Both QNN-1 and QNN-2 are trained using a unified loss function that optimally balances classification loss $C()$ and divergence loss $D()$.
IV-C Methodology Overview
@@ IV-D Training Methodology
Figure 4: Figure depicting the STIQ inference methodology: (a) A user query $X$ is processed by two cloud-hosted QNNs, resulting in obfuscated outputs $y_{1}$ and $y_{2}$. Locally, STIQ aggregates these outputs to produce the correct result $y$ for the user. (b) Illustration of an example explaining how individual models may yield erroneous predictions, yet their combined output through STIQ accurately produces the correct output vector.
@@ IV-E Inference Methodology
Figure 5: Plot demonstrating the test accuracies of different 4-qubit QNNs trained on the Fashion-4 dataset with varying penalties ($\lambda$). It is clear that while increasing the penalty enhances obfuscation, it also adversely affects the combined performance.
V Evaluation
V-A Setup
@@
Device: In our study, we conducted all noiseless experiments on Pennylane’s [30] “lightning.qubit” device using the ”Adjoint” differentiation method to calculate the gradients. This method is known for its efficiency in simulating quantum circuits without noise. For our noisy inferencing, we used “qiskit.aer” device. Given the long queue times and limited availability of real quantum devices and our goal to closely replicate the conditions of these noisy quantum devices, we opted to directly integrate noise models from actual quantum hardware. Specifically, we employed noise models from IBM_Sherbrooke, IBM_Cairo and IBM_Hanoi and incorporated them to the ”qiskit.aer” device to simulate the noise characteristics of these machines accurately. Considering that traditional methods for calculating gradients, such as backpropagation, are not viable on actual quantum devices, we utilized the Simultaneous Perturbation Stochastic Approximation (SPSA) method for gradient calculation in our QNN training. This approach produces noisy gradients, which are more indicative of the performance one would expect from real quantum hardware, thereby ensuring our experimental results are more aligned with practical quantum computing environments. We also conducted experiments on real quantum hardwares, specifically IBM_Sherbrooke, Brisbane and Kyoto.
@@
TABLE I: Comparison of final test accuracy after training a 4-qubit QNN on Fashion-4 with varying penalties.
@@
Training: For evaluation, we used PQC-4 and 19 from [31] to build our QNNs (PQC-$x$ represents circuit-$x$ in [31]), initialized with random weights. In most of the experiments we used a 4-qubit or 8-qubit QNN, consistent with other QML works [32, 33]. The baseline QNN utilizes a generic training procedure that relies solely on the classification loss (i.e. CrossEntropyLoss()) to perform the required classification tasks. In contrast, QNN-1 and QNN-2 are trained concurrently using our STIQ methodology, which incorporates a combination of both classification and divergence losses, as explained in Section IV-D. For embedding classical features to their corresponding quantum state we use the angle encoding technique with $RZ$ gate and for measurement, we calculate the Pauli-Z basis expectation value over all the qubits. For training the QNNs we used CrossEntropyLoss() as our classification loss function and CosineSimilarity() as the divergence loss function (if not mentioned otherwise). The hyperparameters used for training are; Epochs: 30, learning_rate ($\eta$): 0.01, batch_size = 32 and optimizer: Adam. All training are done on an Intel Core-i7-12700H CPU with 40GB of RAM.
Dataset: Since NISQ devices struggle with large images due to limited qubits, high error rates and complex data encoding, similar to recent works [32, 34], we conduct all experiments using a reduced feature set of MNIST, Fashion, Kuzushiji and Letters datasets with latent dimension $d=8$ (from original 28$\times$28 image) generated using a convolutional autoencoder [35]. Thus, for each dataset, we create a smaller 4-class dataset from these reduced feature sets i.e., MNIST-4 (class 0, 1, 2, 3), Fashion-4 (class 6, 7, 8, 9), Kuzushiji-4 (class 3, 5, 6, 9) and Letters-4 (class 1, 2, 3, 4) with each having 1000 samples (700 for training and 300 for testing). Since we use 4-qubit QNN models for training and each of these datasets is of dimension $d=8$, we encode 2 features per qubit. Number of shots/trials is set to 1000 for all experiments.
TABLE II: Final test accuracies after training a QNN on MNIST-4 with different divergence loss functions.
@@
Figure 6: Test accuracies of 4-qubit QNNs trained on the MNIST-4 dataset with various divergence loss functions ($L_{D}$).
TABLE III: Comparison of final test accuracy and loss after training a 8-qubit 2-layer QNN on different datasets for 30 epochs with learning rate $0.01$ and penalty ($\lambda$): 0.3. We can see that eventhough performance on individual QNNs are greatly obfuscated, performance of their combined outputs is similar to that of the baseline model.